OpenStudy (anonymous):

Solve for x: 2x^3=686

OpenStudy (anonymous):

\[x^3=\frac{686}{2}=343\] \[x=\sqrt[3]343=7\]

OpenStudy (anonymous):

Divide both sides by 2 x^3=343 Take the cube root of both sides x = 7

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

On behalf of satellite and myself, it is a sincere pleasure{)